OpenStudy (anonymous):

Ofelia has a certain amount of money. If she spends $28 then she has 3/5 of the original amount left. How much money did she have originally? Write an equation and solve.

OpenStudy (jdoe0001):

so if she spends 28, she's left with only 3/5 of the original amount 2/5 + 3/5 = 5/5 or the whole amount so 28 is really 2/5 of the whole so we could say \(\large { \begin{array}{ccllll} amount&ration \\\hline\\ 28&\frac{2}{5}\\ x&\frac{1}{1} \end{array}\implies \cfrac{\frac{2}{5}}{\frac{1}{{\color{blue}{ 1}}}}\implies \cfrac{28}{x}=\cfrac{2}{5}\cdot \cfrac{{\color{blue}{ 1}}}{1} }\) solve for "x"